# Poincare field theory for massless particles

### v2

@@ 11.1 Conditions for highest helicity in coordinate representation satisfy Fierz-Pauli constraints
$=\int d^{4}p\Big{[}\alpha\Big{(}T^{A}_{+n}(p)+T^{A}_{+n}(-p)\Big{)}+\beta\Big{(}T^{A}_{-n}(p)+T^{A}_{-n}(-p)\Big{)}\Big{]}e^{-ipx}\,.$ (1)
It satisfy massless Klein-Gordon equation and has following gauge transformation
@@
when parameter satisfy Fronsdal condition [7]
@@
11.2 Derivation of Fronsdal equations
@@
This shows that field equations consists of massless Klein-Gordon equation plus supplementary conditions. Supplementary conditions define the minimal set of auxiliary fields suitable for the description of the action found in Refs.[6] and [7].
Above equations can be combined into one equation as linear combination of massless Klein-Gordon equation and supplementary conditions
$\partial^{2}h_{A}+\sum_{i=1}^{n}A_{a_{i}}\partial^{b}h_{bA_{i}}+\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}B_{a_{i}}C_{a_{j}}h^{b}{}_{bA_{ij}}=0\,.\qquad$ (6)
All coefficients can depend on derivatives $\partial_{a_{i}}$ so that we can chose $A_{a_{i}}=a_{i}\partial_{a_{i}},B_{a_{i}}=b_{i}\partial_{a_{i}}$, $C_{a_{i}}=c_{i}\partial_{a_{i}}$ and obtain
$\partial^{2}h_{A}+\sum_{i=1}^{n}a_{i}\partial_{a_{i}}\partial^{b}h_{bA_{i}}+\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}a_{i}b_{j}\partial_{a_{i}}\partial_{a_{j}}h^{b}{}_{bA_{ij}}=0\,.\qquad$ (7)
We can find coefficients $a_{i}$, $b_{i}$ and $c_{i}$ from requirements that this equation is gauge invariant. So, we will require that
$\partial^{2}\delta h_{A}+\sum_{i=1}^{n}a_{i}\partial_{a_{i}}\partial^{b}\delta h_{bA_{i}}+\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}a_{i}b_{j}\partial_{a_{i}}\partial_{a_{j}}\delta h^{b}{}_{bA_{ij}}=0\,.\qquad$ (8)
Using expression (12) and (3) we can find particular gauge transformations
@@
$\delta h^{b}{}_{bA_{ij}}=2\partial^{b}\Omega_{bA_{ij}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}=2\partial^{b}\Omega_{bA_{ij}}\,.\qquad$ (9)
The field equation is gauge invariant if
$\partial^{2}\sum_{i=1}^{n}(1+a_{i})\partial_{a_{i}}\Omega_{A_{i}}+\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}(2b_{i}c_{j}+a_{i})\partial_{a_{i}}\partial_{a_{j}}\partial^{b}\Omega_{bA_{ij}}=0\,.\qquad$ (10)
It produces $a_{i}=-1$ and $b_{i}=c_{i}=\frac{1}{\sqrt{2}}$ so that we obtain Fronsdal equations [7]
$R_{A}(h)\equiv\partial^{2}h_{A}-\sum_{i=1}^{n}\partial_{a_{i}}\partial^{b}h_{bA_{i}}+\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}\partial_{a_{i}}\partial_{a_{j}}h^{b}{}_{bA_{ij}}=0\,,\qquad$ (11)
which is gauge invariant for traceless gauge parameter.
@@ 11.2.1 Fronsdal action
If we consider (11) as Ricci like tensor, then scalar like tensor is
$R_{A_{ij}}=\eta^{a_{i}a_{j}}R_{A}=2\partial^{2}h^{a}{}_{aA_{ij}}-2\partial^{a}\partial^{b}h_{abA_{ij}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\partial^{b}h^{a}{}_{abA_{ijk}}\,,\qquad\qquad$ (12)

### v3

@@ 11 Relations with previous work
$=\int d^{4}p\Big{[}\alpha\Big{(}T^{A}_{+n}(p)+T^{A}_{+n}(-p)\Big{)}+\beta\Big{(}T^{A}_{-n}(p)+T^{A}_{-n}(-p)\Big{)}\Big{]}e^{-ipx}\,,$ (1)
and gauge parameters $\Omega^{A_{i}}(x)$ in coordinate representation
$\Omega^{A_{i}}(x)=\int d^{4}p\Big{(}\Omega^{A_{i}}(p)+\Omega_{A_{i}}(-p)\Big{)}e^{-ipx}$
$=\int d^{4}p\Big{[}\alpha\Big{(}\Omega^{A_{i}}_{+n}(p)+\Omega^{A_{i}}_{+n}(-p)\Big{)}+\beta\Big{(}\Omega^{A_{i}}_{-n}(p)+\Omega^{A_{i}}_{-n}(-p)\Big{)}\Big{]}e^{-ipx}\,.$ (2)
The fields satisfy massless Klein-Gordon equation and has following gauge transformation
@@
when gauge parameters according to (13) satisfy Fronsdal condition [7]
@@
Later we will see that this is necessary condition that equation of motion is gauge invariant.
@@
11.2 Fronsdal approach
In this subsection we will derive Fronsdal equations and Fronsdal action [7]. The corresponding field consist of two components with highest helicity, which are one dimensional. Consequently the field has two degrees of freedom. To satisfy that requirement we must impose a condition on the field to be double-traceless.
11.2.1 Derivation of Fronsdal equations
@@
This shows that field equations consists of massless Klein-Gordon equation plus supplementary conditions (gauge fixing). Supplementary conditions define the minimal set of auxiliary fields suitable for the description of the action found in Refs.[6] and [7].
Let us combine two conditions from (6) into one condition known as de Donder gauge
$G_{A_{i}}=\partial^{b}h_{bA_{i}}-\frac{1}{2}\sum_{j=1\,j\neq i}^{n}\partial_{a_{j}}h^{b}{}_{bA_{ij}}=0\,.\qquad$ (7)
Using expression (12) we can find particular gauge transformations
@@
$\delta h^{b}{}_{bA_{ij}}=2\partial^{b}\Omega_{bA_{ij}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}=2\partial^{b}\Omega_{bA_{ij}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}\,.\qquad$ (8)
Then gauge variation of the de Donder tensor is
$\delta G_{A_{i}}=\partial^{b}\Big{(}\partial_{b}\Omega_{A_{i}}+\sum_{j=1\,j\neq i}^{n}\partial_{a_{j}}\Omega_{bA_{ij}}\Big{)}-\frac{1}{2}\sum_{j=1\,j\neq i}^{n}\partial_{a_{j}}\Big{(}2\partial^{b}\Omega_{bA_{ij}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}\Big{)}$
$=\partial^{2}\Omega_{A_{i}}-\frac{1}{2}\sum_{j=1\,j\neq i}^{n}\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{j}}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}\,.\qquad$ (9)
Equations (6) can be combined into one equation as linear combination of massless Klein-Gordon equation and supplementary conditions in terms of de Donder gauge
$\partial^{2}h_{A}+\sum_{i=1}^{n}a_{i}\partial_{a_{i}}G_{A_{i}}=0\,.\qquad$ (10)
We can find coefficients $a_{i}$, from requirements that this equation is gauge invariant which means that
$\partial^{2}\sum_{i=1}^{n}\partial_{a_{i}}\Omega_{A_{i}}+\sum_{i=1}^{n}a_{i}\partial_{a_{i}}\Big{(}\partial^{2}\Omega_{A_{i}}-\frac{1}{2}\sum_{j=1\,j\neq i}^{n}\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{j}}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}\Big{)}=0\,.\qquad$ (11)
This is satisfied for $a_{i}=-1$ and
$\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{i}}\partial_{a_{j}}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}=0\,.$ (12)
Consequently, in order to ensure that the Fronsdal equation is gauge invariant, we must require that the gauge parameter is traceless
$\Omega^{b}{}_{bA_{ijk}}=0\,.$ (13)
In our approach, this condition is satisfied, as we have shown in (4).
So, equation of motion takes the form of Ricci like tensor
$R_{A}(h)\equiv\partial^{2}h_{A}-\sum_{i=1}^{n}\partial_{a_{i}}G_{A_{i}}=\partial^{2}h_{A}-\sum_{i=1}^{n}\partial_{a_{i}}\partial^{b}h_{bA_{i}}+\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}\partial_{a_{i}}\partial_{a_{j}}h^{b}{}_{bA_{ij}}=0\,.\qquad$ (14)
@@ 11.2.2 Fronsdal action
Then scalar like tensor is
$R_{A_{ij}}=\eta^{a_{i}a_{j}}R_{A}=\partial^{2}h^{a}{}_{aA_{ij}}-2\partial^{a}G_{aA_{ij}}-\sum_{i=1}^{n}\partial_{a_{i}}G^{a}{}_{aA_{ijk}}$
$=2\partial^{2}h^{a}{}_{aA_{ij}}-2\partial^{a}\partial^{b}h_{abA_{ij}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\partial^{b}h^{a}{}_{abA_{ijk}}\,,\qquad\qquad$ (15)
@@
11.2.3 Propagating degrees of freedom of Fronsdal field
From general consideration we know that highest helicity representation is one dimensional and carry one degree of freedom. Then field in Fronsdal equation consists of two highest helicity components and must carry two degrees of freedom. We will assume that it is $k$-traceless tensor and we will find $k$ so that corresponding tensor carries two degrees of freedom. Note that the condition $n\geq 2k$ must be satisfied.
Completely symmetric $k$-traceless tensor with helicity $n$ has
$h^{k}_{A}={n+3\choose 3}-{n+3-2k\choose 3}\,,\qquad\qquad$ (19)
independent components. After sme calculation we obtain
$h^{k}_{A}=n^{2}k-2nk(k-2)-4k^{2}+\frac{k}{3}\Big{(}4k^{2}+11\Big{)}\,.\qquad\qquad$ (20)
Some of these components are related by the gauge condition $G_{A_{i}}=0$. Since de Donder tensor is traceless, $\eta^{ab}G_{abA_{ijk}}=0$, we should imposes
$\Omega^{\prime}_{A_{i}}={n+2\choose 3}-{n\choose 3}=n^{2}\,,\qquad(n\geq 3)\qquad$ (21)
conditions. Note that parameter $\Omega_{A_{i}}$ has one index less then field $h_{A}$. Consequently comparing with (19) in the first term we will take $n\to n-1$. Because parameter $\Omega_{A_{i}}$ is traceless in the second term we will take $k=1$ and $n\to n-1$.
In de Donder gauge, the Fronsdal equation becomes
$\partial^{2}h_{A}=0\,.\qquad$ (22)
Note that de Donder gauge does not fix the gauge completely. Both Fronsdal equation and de Donder gauge are invariant under residual gauge transformation given by traceless gauge parameters $\omega_{A_{i}}$ which obey $\partial^{2}\omega_{A_{i}}=0$. This residual gauge kills additional $\omega^{\prime}_{A_{i}}=n^{2}$ components.
Therefore, since complete number of degrees of freedom must be equal to two we have
$h^{k}_{A}-\Omega^{\prime}_{A_{i}}-\omega^{\prime}_{A_{i}}=n^{2}k-2nk(k-2)-4k^{2}+\frac{k}{3}\Big{(}4k^{2}+11\Big{)}-2n^{2}=2\,.\qquad\qquad$ (23)
This is cubic equation for $k$ with one solution $k=2$. There is one more solution with integers $k=3$ and $n=3$. But it does not satisfy condition $n\geq 2k$.
This means that double-traceless fields for arbitrary helicity in Fronsdal equation indeed propagates two degrees of freedom.
@@ 11.3.1 Generalized curvatures in coordinate representation
12 Arbitrary rank tensor with helicity $\lambda=1$ produces Maxwell equations and with $\lambda=2$ Einstein equations
Arbitrary rank tensor with $n\geq 2$ components contains also helicities $\lambda=\pm 1$ and $\lambda=\pm 2$. We will show that these tensors describe Maxwell equations and Einstein equations in weak field approximation.
12.1 Arbitrary rank tensor with helicities $\lambda=\pm 1$
The components of $n$ rank tensor with helicities $\lambda=\pm 1$ have a form
$T_{\pm 1}^{A}(k)=(P_{\pm 1})^{A}{}_{B}T^{B}(k)=\sum_{i=1}^{n}(\pi_{\pm})^{a_{i}}{}_{b_{i}}(\pi_{0})^{A_{i}}{}_{B_{i}}T^{b_{i}B_{i}}(k)\,.$ (1)
Since $\pi_{0}$ project to two dimensional space and $\pi_{+}$ and $\pi_{-}$ to one dimensional space, both components $T_{\pm 1}^{A}(k)$ are $2^{n-1}n$ dimensional representations and they have $2^{n-1}n$ degrees of freedom.
$T_{\pm 1(-)}^{A}(k)=(P_{\pm 1(-)})^{A}{}_{B}T^{B}(k)=\sum_{i=1}^{n}(\pi_{\pm})^{a_{i}}{}_{b_{i}}(\pi_{0-})^{A_{i}}{}_{B_{i}}T^{b_{i}B_{i}}(k)\,.$ (2)
Note that now each component $T_{+1(-)}^{A}(k)$ and $T_{-1(-)}^{A}(k)$ has $n$ degrees of freedom.
We will consider tensors symmetric in all indices. The other $n-1$ possibilities are antisymmetric expression in $a_{i}$ and one of $n-1$ indices from $A_{i}$. It means that each component $(T)_{+1(-)}^{A}(k)$ and $(T)_{-1(-)}^{A}(k)$ has one degree of freedom. From now on, we will assume that all tensors are totaly symmetric in all indices, but we will not denote symmetrization explicitly. For example $T_{\pm 1(-)}^{a_{i}A_{i}}$ means totaly symmetric in all indices $A$.
Using expressions for projectors we get
$T_{\pm 1(-)}^{a_{i}A_{i}}(k)=-\frac{1}{2^{n}}\Big{(}(\breve{p}_{\mp})_{b_{j}}q_{B_{j}}T^{b_{i}B_{i}}(k)\Big{)}\sum_{i=1}^{n}k^{A_{i}}\breve{p}_{\pm}^{a_{i}}\equiv\mathbf{T}_{\mp n}\sum_{i=1}^{n}k^{A_{i}}\breve{p}_{\pm}^{a_{i}}\,.$ (3)
The fields $T_{\pm 1(-)}^{A}(k)$ with helicities $\lambda=\pm 1$ contain one basic vector $\breve{p}_{\pm}^{a}$ and consequently relation (8) is satisfy.
Therefore, we can write totally symmetric tensor in the form
$(T_{\pm 1(-)})^{A}(k)=\sum_{i=1}^{n}k^{A_{i}}A_{\pm}^{a_{i}}(k)\,,\qquad A_{\pm}^{a}=\mathbf{T}_{\mp n}\breve{p}_{\pm}^{a}\,,$ (4)
Here $A_{\pm}^{a}$ is new field. Note that it is proportional to $\breve{p}_{\pm}^{a}$, as well as fields $V_{\pm}^{a}$ in Eq.(4.3). According to scalar product of basic vectors $k_{a}\breve{p}_{\pm}^{a}=0$ we can conclude that $k_{a}A_{\pm}^{a}=0$, which means that these solutions are in Lorentz gauge. We will see that these fields describe electromagnetic interaction.
12.2 Gauge transformation for arbitrary rank tensor with helicities $\lambda=\pm 1$ and Maxwell equations
Using gauge transformations for basic vectors $\delta k^{a}=0$ and $\delta\breve{p}_{\pm}^{a}=k^{a}\varepsilon_{\pm}$ we can obtain gauge transformations for totally symmetric arbitrary rank tensor with helicities $\lambda=\pm 1$
$\delta(T_{\pm 1(-)})^{A}(k)=k^{A}\omega_{\pm}\,,\qquad\omega_{\pm}(k)=n\mathbf{T}_{\mp n}\varepsilon_{\pm}\,.\qquad$ (5)
and for field $A_{\pm}^{a}$
$\delta A_{\pm}^{a}(k)=k^{a}\tilde{\omega}_{\pm}\,,\qquad\tilde{\omega}_{\pm}(k)=\mathbf{T}_{\mp n}\varepsilon_{\pm}\,.\qquad$ (6)
Since $k^{a}$ and $q^{a}$ are invariant under space inversion and ${\cal P}\breve{p}_{\pm}^{a}=-\breve{p}_{\mp}^{a}$ and ${\cal P}\varepsilon_{\pm}=-\varepsilon_{\mp}$ we have ${\cal P}(T_{\pm 1(-)})^{A}(k)=(T_{\mp 1(-)})^{A}(k)$, ${\cal P}A^{a}_{\pm}(k)=A^{a}_{\mp}(k)$ and ${\cal P}\omega_{\pm}(k)=\omega_{\mp}(k)$. So, according to (11) instead of two components $(T_{+1(-)})^{A}(k)$ and $(T_{-1(-)})^{A}(k)$ we will introduce one symmetric tensor
$T^{A}(k)=\alpha(T_{+1(-)})^{A}(k)+\beta(T_{-1(-)})^{A}(k)\,,\qquad$ (7)
and instead of two components $A^{a}_{\pm}$ we will introduce one vector field
$A^{a}(k)=\alpha A^{a}_{+}(k)+\beta A^{a}_{-}(k)\,,\qquad$ (8)
which are incomplete irreducible representation of Poincare group with space inversion. The representation of the field $T^{A}(k)$ as well as of $A^{a}(k)$ is two dimensional and consequently both fields have two degrees of freedom. Since the parameters $\omega_{\pm}(k)$ and ${\tilde{\omega}}_{\pm}(k)$ are also connected by parity transformation we can introduce one parameter $\omega(k)=\alpha\omega_{+}(k)+\beta\omega_{-}(k)$ and one parameter ${\tilde{\omega}}(k)=\alpha{\tilde{\omega}}_{+}(k)+\beta{\tilde{\omega}}_{-}(k)$
If we go to arbitrary frame and coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$, we obtain expression for totally symmetric tensor $T^{A}(x)$ in terms of vector field $A^{a}(x)$ and theirs gauge transformations. For real field, according to (7) and (8)
$T^{A}(x)=\sum_{i=1}^{n}\partial^{A_{i}}A^{a_{i}}(x)\,.\qquad$ (9)
To obtain gauge transformation we use (7.2) and (6) and obtain
$\delta T^{A}(x)=\partial^{A}\omega(x)\,.\qquad\delta A^{a}(x)=\partial^{a}\omega_{1}(x)\,,\qquad\omega_{1}(x)=\frac{1}{n}\omega(x)\,.$ (10)
The field $A^{a}(x)$ behaves like electromagnetic field. It is vector field, it has the same helicity, the same gauge transformation and the same number of degrees of freedom. So, this describes electromagnetic interaction.
We already mentioned that there are $n-1$ possibe antisymmetric expressions in $a_{i}$ and one of $n-1$ indices from $A_{i}$. It means that each component $(T^{An})_{+1(-)}^{A}(k)$ and $(T^{An})_{-1(-)}^{A}(k)$ has one degree of freedom. Using expressions for projectors we get
$(T^{An})_{\pm 1(-)}^{a_{i}a_{j}A_{ij}}(k)=-\frac{1}{2^{n}}\Big{(}(\breve{p}_{\mp})_{b_{k}}k_{b_{q}}q_{B_{kq}}T^{b_{k}b_{q}B_{kq}}(k)\Big{)}\sum_{i,j=1}^{n}k^{A_{ij}}\Big{(}k^{a_{i}}\breve{p}_{\pm}^{a_{j}}-k^{a_{j}}\breve{p}_{\pm}^{a_{i}}\Big{)}$
$\equiv\mathbf{T}_{\mp n}\sum_{i,j=1}^{n}k^{A_{ij}}\Big{(}k^{a_{i}}\breve{p}_{\pm}^{a_{j}}-k^{a_{j}}\breve{p}_{\pm}^{a_{i}}\Big{)}\,.\qquad$ (11)
Since $k_{a}k^{a}=0$ and $k_{a}\breve{p}_{\pm}^{a}=0$ there are constraints $k_{a_{i}}(T^{An})_{\pm 1(-)}^{a_{i}a_{j}A_{ij}}(k)=0$ and $k_{A_{ij}}T_{(}T^{An})_{\pm 1(-)}^{a_{i}a_{j}A_{ij}}(k)=0$. The fields $(T^{An})_{\pm 1(-)}^{a_{i}a_{j}A_{ij}}(k)$ with helicities $\lambda=\pm 1$ contain one basic vector $\breve{p}_{\pm}^{a}$ and consequently relation (8) is satisfy.
Both $k^{a}$ and $\Big{(}k^{a_{i}}\breve{p}_{\pm}^{a_{j}}-k^{a_{j}}\breve{p}_{\pm}^{a_{i}}\Big{)}$ are gauge invariant. Consequently fields $(T^{An})_{\pm 1(-)}^{a_{i}a_{j}A_{ij}}(k)$ are Lorentz invariant and we can use them as physical fields. Instead of two components $(T^{An}_{+1(-)})^{A}(k)$ and $(T^{An}_{-1(-)})^{A}(k)$, according to (11) we will introduce one antisymmetric tensor
$F^{A}(k)=\alpha(T^{An}_{+1(-)})^{A}(k)+\beta(T^{An}_{-1(-)})^{A}(k)\,,\qquad$ (12)
which is irreducible representation of Poincare group with space inversion. Similarly, from (12.2.1) we have
$F^{A}(k)=k^{A_{ij}}F^{a_{i}a_{j}}\,,\qquad F^{a_{i}a_{j}}=k^{a_{i}}A^{a_{j}}(k)-k^{a_{j}}A^{a_{i}}(k)\,,\qquad$
$A^{a}(k)=\alpha A_{+}^{a}(k)+\beta A_{-}^{a}(k)\,.\qquad$ (13)
Using scalar product of basic vectors we can conclude that there are two constraints on field $F^{ab}(k)$
$k_{a}F^{ab}(k)=0\,,\qquad\varepsilon_{abcd}k^{b}F^{cd}(k)=0\,.\qquad$ (14)
$F_{ab}(x)=\partial_{a}A_{b}(x)-\partial_{b}A_{a}(x)\,,\qquad$ (16)
which also follows from (12.2.1).
12.3 Symmetric tensor with helicities $\lambda=\pm 2$ and free Einstein equations
Similarlly as in the previous subsections we can consider fields with two projectors $(\pi_{\pm})^{a}{}_{b}$ and $n-2$ projectors $(\pi_{0-})^{a}{}_{b}$ (gauge invariant part of $(\pi_{0})^{a}{}_{b}$).
Using expressions for projectors we obtain
$(T^{Sy})_{\pm 2(-)}^{a_{i}a_{j}A_{ij}}(k)=\frac{1}{2^{n}}\Big{(}(\breve{p}_{\mp})_{b_{k}}(\breve{p}_{\mp})_{b_{q}}q_{B_{kq}}T^{b_{k}b_{q}B_{kq}}(k)\Big{)}\sum_{i,j=1}^{n}k^{A_{ij}}\breve{p}_{\pm}^{a_{i}}\breve{p}_{\pm}^{a_{j}}\equiv\mathbf{T}_{\mp n}\sum_{i,j=1}^{n}k^{A_{ij}}\breve{p}_{\pm}^{a_{i}}\breve{p}_{\pm}^{a_{j}}\,.\qquad$ (17)
It is proportional to the tensor (1)
$T_{\pm n}^{ab}(k)=\frac{1}{4}(\breve{p}_{\mp})_{c}(\breve{p}_{\mp})_{d}T^{cd}(k)\breve{p}_{\pm}^{a}\breve{p}_{\pm}^{b}\,,\qquad$ (18)
so that
$(T^{Sy})_{\pm 2(-)}^{a_{i}a_{j}A_{ij}}(k)\sim\sum_{i,j=1}^{n}k^{A_{ij}}T_{\pm 2}^{a_{i}a_{j}}(k)\,.\qquad$ (19)
$H^{a_{i}a_{j}A_{ij}}(x)\sim\sum_{i,j=1}^{n}\partial^{A_{ij}}h^{a_{i}a_{j}}(x)\,.\qquad$ (20)
Therefore, corresponding field is $n-2$ derivative of the field $h^{ab}$ which, as we showed earlier, describe Einstein equations in weak field approximation.